OpenStudy (anonymous):

Evaluate each logarithm. Log(subscript2)16 The answer is 4

OpenStudy (anonymous):

that is because \[log_2(16)\] means the number you raise 2 to to get 16 and\[2^4=16\]

OpenStudy (anonymous):

\[log_2(x)=y\] same as \[2^y=x\]

OpenStudy (anonymous):

Thank you! Now I get it!! :D

OpenStudy (anonymous):

wait what about \[\log_{81}9 \]?

OpenStudy (anonymous):

81 to what power is 9? sqrt(81)=9 so 81^1/2=9 so our answer is 1/2

OpenStudy (anonymous):

OH. thanksssssss!